# Vision LLMs Are Bad at Hierarchical Visual Understanding, and LLMs Are the Bottleneck

### v1

@@ Abstract
This paper reveals that many state-of-the-art large language models (LLMs) lack hierarchical knowledge about our visual world, unaware of even well-established biology taxonomies. This shortcoming makes LLMs a bottleneck for vision LLMs’ hierarchical visual understanding (e.g., recognizing Anemone Fish but not Vertebrate). We arrive at these findings using about one million four-choice visual question answering (VQA) tasks constructed from six taxonomies and four image datasets. Interestingly, finetuning a vision LLM using our VQA tasks reaffirms LLMs’ bottleneck effect to some extent because the VQA tasks improve the LLM’s hierarchical consistency more than the vision LLM’s. We conjecture that one cannot make vision LLMs understand visual concepts fully hierarchical until LLMs possess corresponding taxonomy knowledge.
footnotetext: Code and project page: https://yuanqing-ai.github.io/llm-hierarchy/.
@@ 1 Introduction
Taxonomy is natural and core in visual understanding. The biology taxonomies cover many objects in our visual world [53]; for example, a Boston Terrier belongs to the class of Terrier, which is a subtype of Dog, under Mammal, and ultimately part of the broader category Animal, forming a semantic path in the animal taxonomy: Animal $\rightarrow$ Mammal $\rightarrow$ Dog $\rightarrow$ Terrier $\rightarrow$ Boston Terrier. ImageNet [13] expands from the WordNet [37] taxonomy. Visual parts [28, 15, 3], attributes [14, 27, 41], and relationships [26] can be grouped hierarchically due to shared characteristics.
A high-performing, general-purpose visual understanding system should map visual inputs to both fine-grained leaf nodes of a taxonomy and coarse-grained inner nodes. Meanwhile, it should label an input hierarchically consistently along the path that traces a leaf up to the root. Figure 1 illustrates a case selected from our experiments that the model predictions lack hierarchical consistency, failing to follow the path of Animal $\rightarrow$ Vertebrate $\rightarrow$ Fish $\rightarrow$ Spiny-finned Fish $\rightarrow$ Anemone Fish.
Surprisingly, little has been done to assess the hierarchical visual understanding performance of vision large language models (VLLMs) [4, 29, 9, 72, 34, 29], which have the potential to make such a general-purpose vision system. Indeed, VLLMs unify various vision tasks (e.g., visual recognition [13], captioning [8], question answering [2], and retrieval [62]) into one model by anchoring visual encoders [46, 66, 10, 39] to a versatile pretrained LLM [19, 60], typically orders of magnitude bigger, offering integrated interactions with humans that involve images and videos in conjunction with natural language prompts. Comprehensively benchmarking VLLMs is essential for realizing their potential and identifying opportunities for improvements. Extensive benchmarks have recently emerged, such as the bilingual MMBench [36], manually labeled MME [16], and MMMU [64] collected from college exams. We refer readers to [67] for an extensive list.
This work systematically evaluates VLLMs’ hierarchical visual understanding capabilities using six taxonomies and four hierarchical image classification datasets. Conventionally, the hierarchical image classification [47, 44, 58, 61, 59] aims to classify visual inputs into semantically structured categories across multiple levels of specificity, in contrast to flat classification, which treats labels as mutually exclusive and unstructured. We construct about one million four-choice visual question-answering (VQA) tasks from the hierarchical datasets (see Figure 1 for some examples). The tasks traverse all taxonomy levels, and the four choices of an individual task are from the same level. When evaluating VLLMs’ performance over these tasks, we stress hierarchical consistency because it is unique to hierarchical visual understanding and crucial for adaptability to users’ varying granularity preferences [44, 12, 58].
@@
Figure 1: Left: Four-choice VQA tasks for evaluating VLLMs’ hierarchical visual understanding. Right: A VLLM’s answers (in red boxes) deviate from the ground truth path (green arrows), illustrating its lack of hierarchical consistency.
@@
Our main findings are as follows. First of all, many state-of-the-art VLLMs struggle with our VQA tasks, substantially lacking hierarchical consistency. For example, Qwen2.5-VL-72B [4] makes mistakes over 67% of the hierarchical paths in the iNaturalist [53] taxonomy. Moreover, in our attempt to tracing down the error causes, we find that LLMs are the bottleneck and lack taxonomy knowledge about the visual world. In contrast, the visual encoder and projector modules demonstrate the ability to retain highly discriminative and well-structured visual features. We further show that the LLM embeddings about the visual concepts contain sufficient hierarchical cues and organize them orthogonally, but the model cannot decode them. Finally, finetuning a VLLM using our VQA tasks enhance its LLM’s (text) hierarchical consistency more than the VLLM’s (visual) hierarchical consistency, reaffirming LLMs’ bottleneck effect to some extent.
@@
2 VLLMs Lack Hierarchical Consistency in Visual Understanding
@@
We construct six hierarchical image classification benchmarks in a four-choice VQA format to systematically assess VLLMs’ accuracy and hierarchical consistency in visual understanding. These benchmarks leverage datasets that inherently exhibit taxonomic structures, either derived from WordNet [37] or grounded in biological classification standards. In what follows, we formally define hierarchical image classification, followed by two evaluation metrics about accuracy and consistency, respectively. We then describe our VQA tasks and the first set of experiment results in this work.
2.1 Hierarchical Image Classification: Notations and Problem Statement
@@
General image classification tasks typically assume a flat label space, where each image ${x}\in\mathcal{X}$ is assigned a class label $y\in\mathcal{Y}$ out of a predefined set $\mathcal{Y}$ of mutually exclusive categories. However, many real-world problems exhibit rich semantic structures, in which labels are naturally organized into a hierarchy $\mathcal{T}=(\mathcal{Y},\mathcal{E})$ [44, 58, 61, 59], such as a tree or a directed acyclic graph. Here, $\mathcal{E}\subseteq\mathcal{Y}\times\mathcal{Y}$ denotes the set of directed edges representing parent-child relationships, where $(y_{i},y_{j})\in\mathcal{E}$ indicates that $y_{i}$ is the parent of $y_{j}$ in the hierarchy. In hierarchical image classification, the objective is not only to predict the leaf node label $y\in\mathcal{Y}_{leaf}\subseteq\mathcal{Y}$ but also to correctly recover its full ancestral path $(y_{0},y_{1},\cdots,y_{L})$ in $\mathcal{T}$, where $y_{0}$ denotes the root node and $L$ is the depth of the hierarchy. In this paper, we aim to evaluate VLLMs’ hierarchical image classification capabilities, identify their limitations and underlying causes, and enhance their performance based on these insights.
2.2 Two Evaluation Metrics about Accuracy and Consistency, Respectively
@@
where $N$ is the number of images in the testing set, $L^{i}$ denotes the depth of the hierarchy for the $i$-th input $x^{i}$ and may vary for different tasks in uneven trees, $f_{\theta}:\mathcal{X}\mapsto\mathcal{Y}$ is an image classifier, $\mathcal{Y}_{j}$ represents the set of labels at the $j$-th layer of the hierarchy, and $\mathbbm{1}[\cdot]$ is an indicator function. HCA evaluates whether a model’s predictions are consistent with the entire hierarchical path from the root to a leaf node. Specifically, it measures the proportion of samples for which all ancestor nodes along the predicted paths match the ground truth. This is a stricter metric than flat accuracy and serves as our primary evaluation criterion for hierarchical classification.
Leaf-Level Accuracy $\mathrm{Acc_{leaf}}$ [68, 35, 20]. It cares about the predictions at the most fine-grained level of a taxonomy:
@@
Interestingly, $\mathrm{Acc_{leaf}}$ upper-bounds $\mathrm{HCA}$ because correctly assigning a leaf label $y_{L}$ to an input $x$ contributes to $\mathrm{Acc_{leaf}}$, but it does not increase $\mathrm{HCA}$ unless the model makes no mistake over all nodes in the path $(y_{0},y_{1},\cdots,y_{L})$ connecting the leaf label to the root.
Dataset # Levels # Leaf Nodes # Test Images Hierarchy Distribution
CUB-200-2011 [54] 4 200 5,794 13-37-124-200
iNaturalist-Plant [53] 6 4,271 42.71K 5-14-85-286-1702-4271
iNaturalist-Animal [53] 6 5,388 53.88K 6-27-152-715-2988-5388
ImageNet-Animal [13] 11 397 19.85K 2-10-37-81-123-81-65-41-64-34-2
ImageNet-Artifact [13] 8 491 24.55K 5-40-147-204-162-62-44
Food-101 [5] 4 84 21.00K 6-29-40-24
2.3 VQA Tasks Derived from Hierarchical Image Classification Datasets
VLLMs are the image classifiers $f_{\theta}$ in equations (1) and (2), and one can use language prompts to steer their output to a particular taxonomy level. More concretely, we formalize a VQA task for each image given a desired taxonomy level, $(x^{i},\mathcal{Y}_{j}),i=1,2,\cdots,N,j=1,2,\cdots,L^{i},$ as follows.
VQA Tasks. We derive approximately one million four-choice VQA tasks and six taxonomies from four hierarchical image classification datasets [54, 53, 13, 5] to evaluate VLLMs in a closed-set setting. This setting mitigates the challenge of open-set generation, which involves a prohibitively large prediction space [68] and ambiguous prediction granularity. We test different VQA prompts (provided in Appendix C), and they generally follow this format:
<image> Given the plant in the image, what is its taxonomic classificationat the <hierarchy> (e.g., kingdom) level?A.<similar class> B.<ground truth> C.<similar class> D.<similar class>Answer with the option letter only.(Choices are shuffled in the $\begin{array}[]{l}\text{{<image> Given the plant in the image, what is its taxonomic classification}}\\ \text{{at the <hierarchy> (e.g., kingdom) level?}}\\ \text{{A.<similar class> B.<ground truth> C.<similar class> D.<similar class>}}\\ \text{{Answer with the option letter only.}}\quad\text{(Choices are shuffled in the experiments)}\end{array}$
Arguably, the four-choice VQA tasks are easier than the conventional hierarchical image classification, whose label space is orders of magnitude bigger than four. To compensate this difference, we make sure the four choices are from the same level of a taxonomy and use “confusing labels” in the VQA tasks. Specifically, we use SigLIP [66] to compute the cosine similarity scores between an image and all text labels other than the ground truth (at a particular taxonomy level), selecting the top three most similar labels as the distracting VQA choices. Besides, we provide the results of randomly sampled choices in Appendix B.
Hierarchical Image Classification Datasets. Table 1 summarizes the six taxonomies and four datasets we use to construct the VQA tasks. CUB-200-2011 (CUB-200) [54] is a fine-grained bird dataset containing 200 species. We prompt GPT-4o [1] to map each class to a four-level taxonomy: Order $\rightarrow$ Family $\rightarrow$ Genus $\rightarrow$ Specie. To ensure taxonomic accuracy, we cross-validate the generated hierarchy using corresponding entries from Wikipedia. In addition, we incorporate the iNaturalist-2021 (iNat21) dataset [53], a large-scale collection with species-level annotations spanning various biological taxa. We separate it into two taxonomies, Plant and Animal, comprising 4,271 and 5,388 leaf nodes, respectively, and six levels. Both CUB-200 and iNat21 provide well-established biological taxonomies with even hierarchical depths. To increase structural diversity, we also experiment with ImageNet-1K (ImgNet) [13], whose leaf labels are coarser-grained than iNat21 and CUB-200. ImgNet is built upon the WordNet [37]. We extract two relatively well-structured subsets from ImgNet: ImgNet-Animal and ImgNet-Artifact, following [58]. We further refine these subsets to improve label quality and semantic consistency. Food-101 [5] is about food classification, and its hierarchy is constructed based on the recent work of Liang and Davis [32].
Table 2: The hierarchical consistent accuracy (HCA) and leaf-level accuracy $\mathrm{Acc}_{\mathrm{leaf}}$ of six open-source VLLMs, two CLIP-style models, and the proprietary GPT-4o.
@@ 2.4 Experiments and Findings
We mainly study state-of-the-art open-source VLLMs: The Qwen2.5-VL [4] models of 7B, 32B, and 72B parameters, InternVL2.5-8B [9], InternVL3-8B [72], and LLaVA-OV-7B [29]. Meanwhile, we include the proprietary GPT-4o’s results for reference; in general, GPT-4o slightly outperforms Qwen-2.5-VL-72B, but the main findings below still apply. Finally, we experiment with two CLIP-style [46] models, SigLIP-SO400M [66] and OpenCLIP-L [10], following the experiment protocol in [46] except that the candidate labels for each test image are restricted to the same four choices as fed to VLLMs. Table 2 shows the results about the models’ hierarchical consistency (HCA) and leaf-level accuracy ($\mathrm{Acc_{leaf}}$) on iNat21, ImgNet, and CUB-200. The Food-101 results are in Appendix B to save space in the main text. We draw the following conclusions.
VLLMs Lack Hierarchical Consistency in Visual Understanding. Regardless of the leaf-level accuracy, all open-source VLLMs, CLIP models, and GPT-4o lack hierarchical consistency because their HCA is significantly lower than $\mathrm{Acc_{leaf}}$ (up to 99.3% relatively). The gaps on iNat21-Plant are especially big (e.g., 32.82 vs. 55.00 for Qwen2.5-VL-72B and 35.53 vs. 62.95 for GPT-4o). While one might expect better results on ImgNet, neither open-source VLLMs nor GPT-4o can make their HCA match $\mathrm{Acc_{leaf}}$ — more than 20% decrease for all models, indicating that VLLMs make many mistakes along the paths from the taxonomies’ roots to the leaf nodes even when they are correct over the leaves.
Fine-Grained Visual Recognition Remains Challenging for VLLMs. While VLLMs and CLIP models perform moderately well on ImgNet, they struggle with fine-grained object recognition; on the iNat21 dataset, even the best-performing GPT-4o gives rise to only 63% leaf-level accuracy, far from its 86% on ImgNet. Notably, InternVL2.5 and LLaVA-OV’s results (about 27%) on iNat21 are only slightly above random guess (25%), and the CLIP models are barely on par with random guess. In contrast, a small task-specialized model [23] leads to 61.56% leaf-level accuracy on iNat21, and some models [11, 69] achieve 93% accuracy on CUB-200, outperforming all the general-purpose VLLMs in our experiments. These findings are consistent with the recent work [17, 68, 20, 63] that recognizes the limitation of VLLMs on (fine-grained) image classification.
Scaling Law Works for Hierarchical Visual Understanding. Both hierarchical consistency and leaf-level accuracy improve as the size of the Qwen2.5-VL series of models increases. Moreover, the gap between HCA and $\mathrm{Acc_{leaf}}$ progressively narrows. However, the largest models (Qwen2.5-VL-72B and GPT-4o) are still unsatisfactory in terms of both hierarchical consistency and fine-grained recognition, especially on the iNat21 benchmark.
Qwen2.5-VLs Are Among the Most Powerful Open-Source VLLMs. LLaVA-OV-7B’s hierarchical consistency and leaf-level accuracy are below InternVLs and Qwen2.5-VLs. InternVL3-8B improves upon InternVL2.5-8B, but it is still under par with Qwen2.5-VL-7B.
@@
3 Why Are VLLMs Poor at Hierarchical Image Classification?
We systematically investigate potential causes of VLLMs’ low performance on hierarchical visual understanding. We first extensively study prompt variations in Section 3.1 and reveal that some prompts can lead to marginally better results than the rest, but the results remain generally bad. We then examine VLLMs’ visual encoders and subsequent visual tokens to see whether and where essential visual information is lost when it forwards through VLLMs (Section 3.2). Interestingly, the discriminative cues in the visual tokens are maintained across various stages of the VLLM architectures, leading to about the same hierarchical image classification results immediately after the visual encoder, after the projection to the language token space, and at the very last layer of an LLM. Finally and surprisingly, we find that the generally believed powerful LLMs, even the one with 72B parameters in our experiments, lack basic taxonomy knowledge and are likely responsible for VLLMs’ poor performance on hierarchical visual understanding! (We believe this conclusion is true for open-source VLLMs, but we urge readers not to extrapolate it to proprietary LLMs because we could not probe their intermediate embeddings.)

### v2

# The LLM Bottleneck: Why Open-Source Vision LLMs Struggle with Hierarchical Visual Recognition

@@ Abstract
This paper reveals that many open-source large language models (LLMs) lack hierarchical knowledge about our visual world, unaware of even well-established biology taxonomies. This shortcoming makes LLMs a bottleneck for vision LLMs’ hierarchical visual recognition (e.g., recognizing Anemone Fish but not Vertebrate). We arrive at these findings using about one million four-choice visual question answering (VQA) tasks constructed from six taxonomies and four image datasets. Interestingly, finetuning a vision LLM using our VQA tasks reaffirms LLMs’ bottleneck effect because the VQA tasks improve the LLMs’ hierarchical consistency in text-only tasks more than the vision LLMs’. We believe that one cannot make vision LLMs understand our visual world hierarchically until LLMs possess corresponding taxonomy knowledge.
@@ 1 Introduction
We make a potentially controversial argument that large language models (LLMs) can hinder research progress in computer vision because they lack knowledge of our visual world. Indeed, all vision LLMs (VLLMs) examined in this paper, proprietary or open-source, yield unsatisfying results on hierarchical visual recognition. A careful ablation study reveals that LLMs themselves are the bottleneck for open-source VLLMs in this setting, whereas the proprietary ones cannot be ablated.
@@
Taxonomy is natural and core in visual recognition. The biology taxonomies cover more than two million species [63], most of which are visual; for example, a Boston Terrier belongs to the breed of Terrier, which is a subtype of Dog, under Mammal, and ultimately part of the broader category Animal, forming a semantic path in the animal taxonomy: Animal $\rightarrow$ Mammal $\rightarrow$ Dog $\rightarrow$ Terrier $\rightarrow$ Boston Terrier. ImageNet [14] expands from the WordNet [43] taxonomy, and visual parts [32, 17, 4], attributes [16, 31, 50], and relationships [30] can be grouped hierarchically due to shared characteristics.
@@
Figure 1: Left: Four-choice VQA tasks for evaluating VLLMs’ hierarchical visual recognition. Right: A VLLM’s answers (in red boxes) deviate from the ground truth path (green arrows), illustrating its lack of hierarchical consistency.
A high-performing, general-purpose visual recognition system should map visual inputs to both fine-grained leaf nodes of a taxonomy and coarse-grained inner nodes. Meanwhile, it should label an input hierarchically and consistently along the path that traces a leaf up to the root. Figure˜1 illustrates a case selected from our experiments that the model predictions lack hierarchical consistency, failing to follow the path of Animal $\rightarrow$ Vertebrate $\rightarrow$ Fish $\rightarrow$ Spiny-finned Fish $\rightarrow$ Anemone Fish.
Surprisingly, little has been done to assess the hierarchical visual recognition performance of vision large language models (VLLMs) [6, 11, 83, 39, 33, 5, 70, 2], which have the potential to make such a general-purpose vision system. Indeed, VLLMs unify various vision tasks (e.g., visual recognition [14], captioning [10], question answering [3], and retrieval [74]) into one model by anchoring visual encoders [55, 78, 12, 48] to a versatile pretrained LLM [21, 72, 71], typically orders of magnitude bigger, offering integrated interactions with humans that involve images and videos in conjunction with natural language prompts. Comprehensively benchmarking VLLMs is essential for realizing their potential and identifying opportunities for improvements. Extensive benchmarks have recently emerged, such as the bilingual MMBench [41], manually labeled MME [18], and MMMU [76] collected from college exams. We refer readers to [79] for an extensive list.
This work systematically evaluates VLLMs’ hierarchical visual recognition capabilities using six taxonomies and four datasets. Conventionally, hierarchical visual recognition [56, 52, 68, 73, 69] aims to classify visual inputs into semantically structured categories across multiple levels of specificity, in contrast to the flat classification viewing labels as mutually exclusive and unstructured. We construct about one million four-choice visual question-answering (VQA) tasks from the hierarchical datasets (see Fig.˜1 for some examples). The tasks traverse all taxonomy levels, and the four choices of an individual task are from the same level. When evaluating VLLMs’ performance over these tasks, we stress hierarchical consistency because it is unique to hierarchical visual recognition and crucial for adaptability to users’ varying granularity preferences [52, 13, 68].
@@
Our main findings are as follows. First, both proprietary and open-source VLLMs struggle with our VQA tasks, substantially lacking hierarchical consistency. For example, Qwen2.5-VL-72B [6] makes mistakes over 67% of the hierarchical paths in the iNaturalist [63] taxonomy, and GPT-4o is only marginally better. Moreover, we find that LLMs are the bottleneck and lack taxonomy knowledge about the visual world — we believe this conclusion is true for open-source VLLMs, but we urge readers not to extrapolate it to proprietary LLMs because we could not probe their intermediate embeddings. In contrast, the visual encoder and projector modules demonstrate the ability to retain highly discriminative and well-structured visual features. We further show that the LLM embeddings about the visual concepts contain sufficient hierarchical cues and organize them orthogonally, but the model cannot decode them. Finally, finetuning a VLLM helps but cannot fix the issue. Meanwhile, finetuning enhances LLM’s (text) hierarchical consistency more than the VLLM’s (visual) hierarchical consistency, reaffirming LLMs’ bottleneck effect to some extent.
2 Related Work
Hierarchical Recognition. Hierarchical recognition [56, 29] is vital for understanding the visual world [73, 53, 52] and language concepts [81, 66, 82, 26, 46]. Recent studies have shown that CLIP-style [55] models lack consistency across taxonomic levels [68, 20, 69, 49, 15, 1]. To better encode hierarchy structural priors into visual models, hierarchical loss functions, multi-level supervision, and taxonomy-aligned embeddings [77, 57, 9] are introduced. Another line of research [45, 37, 44, 25, 51, 54] aims to understand how hierarchical structures are encoded within pre-trained LLMs. Appendix E provides more details.
Visual Recognition with VLLMs. While VLLMs have demonstrated strong performance across various domains, their effectiveness in fine-grained and subordinate-level recognition remains suboptimal [80, 40, 75, 19, 24, 42, 58]. Zhang et al. [80] first identified the limitations of current VLLMs in fine-grained image classification. Building on this, Liu et al. [40] assess a broader range of models. Beyond closed-set evaluation [75, 19, 24], Conti et al. [13] benchmark VLLMs’ open-world visual recognition, while Snæbjarnarson et al. [58] propose to evaluate VLLMs’ open-set predictions using a taxonomic similarity rather than exact string matching. However, to the best of our knowledge, no prior work has examined VLLMs under the hierarchical visual recognition context.
@@ 2 Related Work
3 Both Proprietary and Open-Source VLLMs Lack Hierarchical Visual Consistency
@@
We construct six hierarchical visual recognition benchmarks in a four-choice VQA format to systematically assess VLLMs’ accuracy and hierarchical consistency in visual recognition. These benchmarks leverage datasets with inherent taxonomic structures, either derived from WordNet [43] or grounded in biological classification standards.
In what follows, we formally define hierarchical visual recognition, followed by two evaluation metrics about accuracy and consistency, respectively. We then describe our VQA tasks and the first set of experiment results.
3.1 Hierarchical Visual Recognition
@@
General visual recognition tasks typically assume a flat label space, where each image ${x}\in\mathcal{X}$ is assigned a class label $y\in\mathcal{Y}$ out of a predefined set $\mathcal{Y}$ of mutually exclusive categories. However, many real-world problems exhibit rich semantic structures, in which labels are naturally organized into a hierarchy $\mathcal{T}=(\mathcal{Y},\mathcal{E})$ [52, 68, 73, 69], such as a tree or a directed acyclic graph. Here, $\mathcal{E}\subseteq\mathcal{Y}\times\mathcal{Y}$ denotes the set of directed edges representing parent-child relationships, where $(y_{i},y_{j})\in\mathcal{E}$ indicates that $y_{i}$ is the parent of $y_{j}$ in the hierarchy. In hierarchical visual recognition, the objective is not only to predict the leaf node label $y\in\mathcal{Y}_{leaf}\subseteq\mathcal{Y}$ but also to correctly recover its full ancestral path $(y_{0},y_{1},\cdots,y_{L})$ in $\mathcal{T}$, where $y_{0}$ denotes the root node and $L$ is the depth of the hierarchy. In this paper, we aim to evaluate VLLMs’ hierarchical visual recognition capabilities, identify their limitations and underlying causes, and enhance their performance based on these insights.
3.2 Evaluation: Accuracy and Consistency
@@
where $N$ is the number of images in the testing set, $L^{i}$ denotes the depth of the hierarchy for the $i$-th input $x^{i}$ and may vary for different tasks in uneven trees, $f_{\theta}:\mathcal{X}\mapsto\mathcal{Y}$ is an image classifier, $\mathcal{Y}_{j}$ represents the set of labels at the $j$-th level of the hierarchy, and $\mathbbm{1}[\cdot]$ is an indicator function. HCA evaluates whether a model’s predictions are consistent with the entire hierarchical path from the root to a leaf node. Specifically, it measures the proportion of test examples for which all ancestor nodes along the predicted paths match the ground truth. This is a stricter metric than flat accuracy and serves as our primary evaluation criterion for hierarchical classification.
Leaf-Level Accuracy $\mathrm{Acc_{leaf}}$ [80, 40, 24]. It cares about the predictions at the most fine-grained leaf level $\mathcal{Y}_{L}$ of a taxonomy:
@@
Interestingly, $\mathrm{Acc_{leaf}}$ upper-bounds $\mathrm{HCA}$ because correctly assigning a leaf label $y_{L}$ to an input $x$ contributes to $\mathrm{Acc_{leaf}}$, but correct leaf-level prediction does not increase $\mathrm{HCA}$ unless the model makes no mistake over all nodes in the path $(y_{0},y_{1},\cdots,y_{L})$ connecting the leaf node to the root.
Table 1: The hierarchical consistent accuracy (HCA) and leaf-level accuracy $\mathrm{Acc}_{\mathrm{leaf}}$ of ten open-source VLLMs, four CLIP-style models, and the proprietary GPT-4o.
@@
3.3 VQA Tasks Derived from Hierarchical Visual Recognition Datasets
In this work, VLLMs are the image classifiers $f_{\theta}$ in equations (1) and (2), and we use language prompts to steer their output to a particular taxonomy level. Concretely, we formalize a VQA task for each image given a desired taxonomy level, $(x^{i},\mathcal{Y}_{j}),i=1,2,\cdots,N,j=1,2,\cdots,L^{i}.$
VQA Tasks. We derive approximately one million four-choice VQA tasks and six taxonomies from four hierarchical visual recognition datasets [64, 63, 14, 7] to evaluate VLLMs in a closed-set setting. This setting mitigates the challenge of open-set generation, which involves a prohibitively large prediction space [80] and ambiguous prediction granularity. We test different VQA prompts (provided in Appendix C), and they generally follow this format:
[…]
Arguably, the four-choice VQA tasks are easier than the conventional hierarchical visual recognition, whose label space is orders of magnitude bigger than four. To compensate this difference, we make sure the four choices are from the same level of a taxonomy and use “confusing labels” in the VQA tasks. Specifically, we use SigLIP [78] to compute the cosine similarity scores between an image and all text labels other than the ground truth (at a particular taxonomy level), selecting the top three most similar labels as the distracting VQA choices. Besides, we provide the results of randomly sampled choices in Appendix B.
Hierarchical Visual Recognition Datasets. CUB-200-2011 (CUB-200) [64] is a fine-grained bird dataset containing 200 species and each class is mapped to a four-level taxonomy: Order $\rightarrow$ Family $\rightarrow$ Genus $\rightarrow$ Species. In addition, we incorporate the iNaturalist-2021 (iNat21) dataset [63], a large-scale collection with species-level annotations spanning various biological taxa. We separate it into two taxonomies, Plant and Animal, comprising 4,271 and 5,388 leaf nodes, respectively, and six levels. To increase structural diversity, we also experiment with ImageNet-1K (ImgNet) [14], whose leaf labels are coarser-grained than iNat21 and CUB-200. ImgNet is built upon the WordNet [43]. We extract two relatively well-structured subsets from ImgNet: ImgNet-Animal and ImgNet-Artifact, following [68]. We further refine these subsets to improve label quality and semantic consistency. Food-101 [7] is about food classification, and its hierarchy is constructed based on the recent work of Liang and Davis [36]. More details are provided in Appendix A.
@@ 3.4 Experiments and Findings
We mainly study state-of-the-art open-source VLLMs: The Qwen2.5-VL [6] models of 7B, 32B, and 72B parameters, Qwen3-VL [5] models of 8B and 32B parameters, Qwen3-Omni-30B-A3B [70], InternVL2.5-8B [11], InternVL3-8B [83], LLaVA-OV-7B [33], and LLaVA-OV-1.5-8B [2]. Meanwhile, we include the proprietary GPT-4o’s results for reference; in general, GPT-4o slightly outperforms Qwen-2.5-VL-72B, but the main findings below still apply. Besides, we experiment with two CLIP-style [55] general-purpose models, SigLIP-SO400M [78] and OpenCLIP-L [12], following the experiment protocol in [55] except that the candidate labels for each test image are restricted to the same four choices as fed to VLLMs. Finally, we evaluate two domain-specific CLIP-style models, BioCLIP [59] and BioCLIP2 [22] under our experimental setup. Table˜1 shows the results about the models’ hierarchical consistency (HCA) and leaf-level accuracy ($\mathrm{Acc_{leaf}}$) on iNat21, ImgNet, and CUB-200. The Food-101 results are in Appendix B to save space. We draw the following conclusions.
Both Proprietary and Open-Source VLLMs Lack Hierarchical Consistency in Visual recognition. Regardless of the leaf-level accuracy, all open-source VLLMs, CLIP models , and GPT-4o lack hierarchical consistency because their HCA is significantly lower than $\mathrm{Acc_{leaf}}$ (up to 99.3% relatively). The gaps on iNat21-Plant are especially big (e.g., 32.82 vs. 55.00 for Qwen2.5-VL-72B and 35.53 vs. 62.95 for GPT-4o). While one might expect better results on ImgNet, neither open-source VLLMs nor GPT-4o can make their HCA match $\mathrm{Acc_{leaf}}$ — more than 20% decrease for all models, indicating that VLLMs make many mistakes along the paths from the taxonomies’ roots to the leaf nodes even when they are correct over the leaves.
Fine-Grained Visual Recognition Remains Challenging for VLLMs, Proprietary or Open-Source. While the general-purpose VLLMs and CLIP models perform moderately well on ImgNet, they struggle with fine-grained object recognition; on the iNat21 dataset, even the best-performing GPT-4o gives rise to only 63% leaf-level accuracy. Notably, InternVL2.5 and LLaVA-OV’s results (about 27%) on iNat21 are only slightly above random guess (25%), and the general CLIP models are barely on par with random guess. In contrast, the domain specific model, BioCLIP2, yields 95.26% leaf-level accuracy on iNat21 and 92.94% accuracy on CUB-200, outperforming the general-purpose VLLMs. These findings are consistent with the recent work [19, 80, 24, 75] that study the limitation of VLLMs on (fine-grained) visual recognition.
@@
Scaling Law Works for Hierarchical Visual Recognition, Only to Some Degree. Hierarchical consistency and leaf-level accuracy improve as the size of the Qwen2.5-VL series of models increases. The gap between HCA and $\mathrm{Acc_{leaf}}$ progressively narrows. However, the largest models (Qwen2.5-VL-72B and GPT-4o) are still unsatisfactory in terms of both hierarchical consistency and fine-grained recognition, especially on the iNat21 benchmark.
Qwen3-VLs and Qwen2.5-VLs Are Among the Most Powerful Open-Source VLLMs. LLaVA-OV-7B’s and LLaVA-OV-1.5-8B’s hierarchical consistency and leaf-level accuracy are below InternVLs and Qwen-series models. InternVL3-8B improves upon InternVL2.5-8B, but it is still under par with Qwen2.5-VL-7B and Qwen3-VL-8B.
Domain-Specific BioCLIPs Lack Hierarchical Consistency. While the domain-specific BioCLIPs yield superb leaf-level accuracy on the iNat21, they still lack hierarchical consistency; there exist persistent gaps between BioCLIPs’ HCA and leaf-level accuracy on all the datasets. It is also worth noting that BioCLIPs’ performance on ImgNets’ leaf levels is below that of the general CLIP models, likely due to their (overly) specialization in the biological domain.
4 Why Are VLLMs Poor at Hierarchical Visual Recognition?
We systematically investigate potential causes of VLLMs’ low performance on hierarchical visual recognition. We first extensively study prompt variations in Section 4.1 and reveal that some prompts can lead to marginally better results than the rest, but the results remain generally bad. We then examine open-source VLLMs’ visual encoders and subsequent visual tokens to see whether and where essential visual information is lost when it forwards through VLLMs (Section 4.2). Interestingly, the discriminative cues in the visual tokens are maintained across various stages of the VLLM architectures, leading to about the same hierarchical visual recognition results immediately after the visual encoder, after the projection to the language token space, and at the very last layer of an LLM.
Finally and surprisingly, we find that the generally believed powerful LLMs, even the one with 72B parameters in our experiments, lack the taxonomy knowledge of our visual world, hence likely responsible for VLLMs’ poor performance on hierarchical visual recognition! (We believe this conclusion holds for open-source VLLMs, but we urge readers not to extrapolate it to proprietary ones because we could not probe their intermediate embeddings.)